OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write the solution as an ordered pair.

OpenStudy (anonymous):

25=m+5f 40=m+10f

OpenStudy (anonymous):

then 40-10f=m sub into other eqn

OpenStudy (anonymous):

25=(40-10f)+5f 25=40-5f -15=-5f 3=f